OpenStudy (anonymous):

If x, y, and z are positive numbers and 2x=3y=4z, then the value of x+y+z is how many times the value of x?

OpenStudy (anonymous):

a. 13/6 b. 17/6 c. 13/2 d. 17/2 e. cannot be determined

hartnn (hartnn):

from 2x=3y=4z, we get y= (2/3)x and z=(1/2)x so x+y+z=x+(2/3)x+(1/2)x =x(1+2/3+1/2) =x(3/2+2/3) =x(13/6)